OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2)

OpenStudy (smileyxl3):

First, find the slope between the two points. m = (y₂ - y₁)/(x₂ - x₁) m = (3 + 2)/(5 - 8) m = 5/-3 = -5/3. Now, use point slope form, with (8, -2). y - y₁ = m(x - x₁) y + 2 = (-5/3)(x - 8) y + 2 = (-5/3)x + 40/3 y = (-5/3)x + 34/3 (5/3)x + y = 34/3 ← standard form is Ax + By = C, but clear the fractions. Multiply by 3 on both sides. 5x + 3y = 34

OpenStudy (smileyxl3):

@DjNck